Using the variable v to represent Vidya's age, translate the following sentence into a mathematical expression:


40 is 4/7 of Vidya's age


Answer by tutorcecilia(2152)

Let x= Vidya's age
4/7x=40
Solving for x:
(4/7)(7/4)x=(7/4)(40)
x=70